# Weighted Ensemble Models Are Strong Continual Learners

### v1

@@ Abstract
In this work we study the problem of continual learning (CL) where the goal is to learn a model on a sequence of tasks, such that the data from the previous tasks becomes unavailable while learning on the current task data. CL is essentially a balancing act between being able to learn on the new task (i.e plasticity) and maintaining the performance on the previously learned concepts (i.e stability). With an aim to address the stability-plasticity trade-off we propose to perform weight-ensembling of the model parameters of the previous and current task. This weight-ensembled model, which we call Continual Model Averaging (or CoMA), attains high accuracy on the current task by leveraging plasticity, while not deviating too far from the previous weight configuration, ensuring stability. We also propose an improved variant of CoMA, named Continual Fisher-weighted Model Averaging (or CoFiMA), that selectively weighs each parameter in the weight ensemble by leveraging the Fisher information of the weights of the model. Both the variants are conceptually simple, easy to implement and effective in attaining state-of-the-art performance on several standard CL benchmarks. [^1] Code is available: https://github.com/IemProg/CoFiMA
@@ 1 Introduction
Deep learning models have achieved impressive results in many tasks such as image classification [20, 11], object detection [64] and semantic segmentation [27]. These advances are largely due to training with extensive offline datasets [54, 8] for a finite set of categories. Contrary to these controlled conditions, real-world applications often involve continuous learning from data streams that introduce novel categories [19, 17]. This process typically results in catastrophic forgetting (CF) [17], where the acquisition of new information leads to the erosion of the previously learned knowledge. Continual Learning (CL) emerges as a solution, allowing models to assimilate new classes from the temporal data stores while retaining classification capabilities for the previously learned classes using a unified classifier [37].
In prior work on CL, the common approach is to train the networks (usually ResNets [20]) from scratch on the current task while regularizing them to counteract forgetting [37, 70]. The emergence of pre-trained models (PTMs) [13, 11, 50] (i.e Vision Transformer (ViT) [11, 36]), trained on large datasets such as ImageNet [54], has provided an effective alternative for CL tasks [72, 73, 71, 66, 84]. PTM-based CL methods [82, 72, 73] enhance data efficiency and consistently excel over traditional CL methods. In particular, Zhang et al. [82] showed that a simple solution such as sequential fine-tuning using PTMs yields competitive results, outperforming modern CL approaches involving prompting [72, 73] and the traditional ones using regularization [28, 33]. This is further validated in recent works [12, 40], where it has been shown that PTMs, owing to their strong and diverse pre-training, are inherently less susceptible to catastrophic forgetting.
Figure 1: Comparison of CoFiMA with sequential fine-tuning in Continual Learning. In sequential finetuning, the model from the previous task serves as the starting point for the current task. Conversely, CoFiMA employs an averaging strategy: the model from the previous task $t-1$ is merged with the one fine-tuned on the current task $t$, using a Fisher-weighted averaging (FWA) technique.
Despite the significant advancements in CL with PTMs, there remains a challenge in achieving an optimal balance between retaining knowledge from previous tasks and excelling in current tasks, referred to as stability-plasticity dilemma [42, 26, 1]. Current methods often face trade-offs between these objectives, leading to either partial forgetting of old tasks or having sub-optimal performance on new ones. In search of finding this optimal balance, we look at CL through the lens of model averaging [51, 79, 75, 74]. Specifically, WiSE-FT [75], a weight-space ensembling approach, that linearly interpolates between the weights of the PTM (e.g., zero-shot CLIP) and the PTM fine-tuned on the target distribution. As a consequence, WiSE-FT achieves high accuracy on the target distribution, while maintaining good performance on datasets with distributional shifts (a setting where zero-shot CLIP would normally excel).
Motivated by the findings of WiSE-FT, we cast CL as a transfer learning approach where we aim to linearly combine the weights from a previous task with the model weights from the current task (see Fig. 1). This endows us with improved plasticity on the current task, while maintaining stability on past tasks. The underpinning idea of our proposed method, Continual Model Averaging (CoMA), also resonates with the work by Mizradeh et al. [41], that compares minima resulting from multitask learning and continual learning, establishing that minima from continual learning are linearly connected to optimal sequential multitask minima but not to each other, leading to forgetting.
@@
The model averaging in CoMA assumes that all the weights of the network have same importance for a given task. While effective, putting equal importance to all weights could result in a weight-ensembled model that lies in an error basin with high loss [74]. To mitigate this problem, we aim to selectively ensemble the weights based on the importance of each weight parameter for a given task. Inspired by elastic weight consolidation (EWC) [28], we leverage the Fisher information [10, 59] to weigh the model parameters during model averaging. Fisher information inherently captures the importance of each weight parameter on the dataset (or task) the model has been trained on. We call this variant of CoMA as Continual Fisher-weighted Model Averaging (or CoFiMA). While this variant requires storing the Fisher information matrix after training on each task, it is computationally lightweight as it requires a single forward and backward pass on the current task data, and yields state-of-the-art performance surpassing both CoMA and existing PTM-based CL solutions.
Our contributions are summarised as follows. (i) We propose CoMA, a weight-ensemble inspired CL approach, that addresses the challenging task of stability-plasticity trade-off in CL. (ii) We extend CoMA by employing Fisher information to weight the averaging of parameters (i.e CoFiMA). Fisher information weighting enables us to determine which parameters contain important “information” to previous tasks, and which can be changed freely. (iii) We run extensive experiments on CL benchmarks and demonstrate that CoFiMA despite being simple consistently outperforms PTM-based CL solutions.
Figure 2: Illustration of the parameter trajectory and model averaging in the loss landscape. (a) Shows the trajectory where models are re-initialized at the start of each task leading to disparate solutions. (b) Depicts decreasing $\lambda_{t}$: Uniform Averaging without re-initialization, showing the convergence of model parameters towards a solution that balances between tasks. (c) Recent tasks are given more weight, resulting in a solution that remains close to the latest task’s model while considering previous tasks. The color gradient indicates the loss value, with darker shades representing lower loss.
@@ 2 Related Work
Continual Learning with PTM. Until now, the predominant focus in continual learning has been on the sequential training of deep neural networks from scratch, aiming to proficiently acquire new tasks while mitigating significant forgetting of preceding tasks. Noteworthy strategies encompass regularization-based approaches [28, 2, 80, 34, 9], which maintain the initial model and selectively stabilize parameter or prediction alterations; replay-based approaches [69, 77, 49, 4], that seek to approximate and regenerate previously learned data distributions; and architecture-based approaches [78, 56, 55], which allocate discrete parameter sub-spaces for each incrementally introduced task.
@@
The recent trajectory of research has further probed into the advantages of pre-training and using PTMs within the CL framework [73, 72, 12]. Representations derived from supervised pre-training have demonstrated the capacity to facilitate not only knowledge transfer but also resilience against catastrophic forgetting during downstream continual learning [52, 39]. Moreover, learning a substantial base of classes during the initial training phase permits CL with minimal adaptations [76]. L2P [73] leveraged techniques inspired by pre-trained knowledge utilization in NLP, employing an additional set of learnable parameters, termed “prompts”, which dynamically guide a pre-trained representation layer in learning incremental tasks. DualPrompt [72] elaborated on this concept by attaching supplementary prompts to the pre-trained representation layer to facilitate the learning of both task-invariant and task-specific instructions. Though prompt-based approaches have been documented to significantly outperform conventional continual learning baselines, they introduce an extra inference cost. Recently, Zhang et al. [82] show that sequential fine-tuning with small learning rate using PTMs outperforms traditional CL approaches.
@@
Unlike prior approaches in CL that enhance PTMs using prompts [72, 73] or replay-buffers [69, 77, 49, 4], CoFiMA employs a different strategy. It reduces forgetting during training by averaging the weights of parameters from previous models. In contrast to EWC [28], which uses the Fisher Information Matrix (FIM) as a regularization term for L2-transfer between tasks, CoFiMA applies the FIM as a weighting factor to assess the significance of weights for each task.
@@
Output-space/weight-space ensembles. Traditional ensemble methods, or output-space ensembles, combine multiple classifiers’ predictions, often outperforming single models and providing more calibrated uncertainty estimates under distribution shifts [18, 32, 46, 61]. These output-space ensembles, however, demand substantial computational resources at inference. Weight-space ensembles offer a computationally efficient alternative by interpolating between model weights [75, 74, 25, 44, 63]. Wortsman et al. [75] achieved this by interpolating between zero-shot CLIP and fine-tuned model weights, resulting in performance gains on both the fine-tuning task and under-distribution shifts. Matena et al. [38] average models with the same initialization that are fine-tuned on different text classification tasks. We show that this merging approach does not produce good performance in CL.
Our approach differs from [75, 74, 38] as it focuses on sequential finetuning with weight-averaging to be accustomed to the CL setting, we iteratively perform our procedure once per task, using the averaged model at each task as the initialization for the next task.
@@
Linear mode connectivity. Interpolation of neural network weights has been shown to maintain high accuracy across various scenarios, along a shared optimization trajectory [16, 75, 14, 7, 25, 74]. The concept of linear mode connectivity, introduced by Frankle et al. [16], describes a state where accuracy is retained along the linear interpolation between two networks’ weights. Analogously, Neyshabur et al. [43] demonstrates that exists a connection between minima obtained by pre-trained models versus freshly initialized ones. They note that there is no performance barrier between solutions coming from pre-trained models, but there can be a barrier between solutions of different randomly initialized models. Mizradeh et al. [41] investigated linear connectivity in the context of multi-task learning and CL, they show that it exists a linear path solution between two models: one excelling in task A, and the other fine-tuned on both tasks A and B.
These works offer a solid foundation of both theoretical and empirical evidence supporting the efficacy of linear interpolation in model performance. Building on these findings, our work presents a novel solution tailored to CL, an area yet to be explored from this angle.
@@ 3.1 Problem formulation and overview
Continual learning with pre-trained models. We consider a classification model $M_{\theta}(\cdot)\!=\!h(f(\cdot))$, where $f(\cdot)$ is a feature extractor and $h(\cdot)$ is a classification head, both parameterized by a unified set of parameters $\theta$. The model $M_{\theta}$ is initialized with parameters $\theta_{0}$ of a pre-trained model and then trained sequentially on a series of incremental tasks, each represented by the corresponding training set $\mathcal{D}_{t}$, for $t\in\{1,\ldots,T\}$. The primary objective is to achieve robust performance across the test sets associated with these tasks. Specifically, for each task $t$, the dataset $\mathcal{D}_{t}$ is defined as $\mathcal{D}_{t}=\bigcup_{c\in C_{t}}\mathcal{D}_{t}^{c}$ where $\mathcal{D}_{t}^{c}=\{(x_{n}^{c},y_{n}^{c})\}_{n=1}^{N_{c}}$, and $C_{t}$ denotes the set of novel classes introduced in task $t$. Here, $N_{c}$ represents the number of training instances for each class $c$, with $(x_{n}^{c},y_{n}^{c})$ denoting the $n$-th training instance and its corresponding label. In Class-Incremental Learning (CIL), evaluation is conducted across all observed classes without the need for task-index labels [65].
@@
This problem poses two primary challenges: (i) the necessity to adapt the knowledge acquired from the pre-trained model to new tasks; and (ii) the imperative to maintain the model’s comprehensive learning capabilities to avoid forgetting previously acquired knowledge while assimilating new tasks.
Overview. Our method, named Continual Model Averaging (CoMA), tackles the CIL problem. Our approach is based on the idea that averaging the old and updated model weights proves to be an efficient method of balancing the old and new knowledge. In summary, upon completing each task, we use the obtained model as a starting point for the next task. This starting model is then fine-tuned on the data from the novel task. The task concludes with a weighted average of the initial and updated model parameters, yielding a model that performs effectively on both the past and current tasks. The averaging procedure is detailed in Sec. 3.2.
Furthermore, we introduce Continual Fisher-weighted Model Averaging (CoFiMA), wherein each parameter is set to its weighted average, derived from both the preceding model and the fine-tuned variant. Unlike CoMA, the weighting for each parameter is determined by its Fisher information [59, 10] to provide control over the importance assigned to each weight of the model. This procedure is detailed in Sec. 3.3, and the corresponding algorithm is described in Alg. 1.
3.2 Continual Model Averaging

### v3

@@ Abstract
In this work, we study the problem of continual learning (CL) where the goal is to learn a model on a sequence of tasks, under the assumption that the data from the previous tasks becomes unavailable while learning on the current task data. CL is essentially a balancing act between learning on the new task (i.e. plasticity) and maintaining the performance on the previously learned concepts (i.e. stability). To address the stability-plasticity trade-off, we propose to perform weight-ensembling of the model parameters of the previous and current tasks. This weighted-ensembled model, which we call Continual Model Averaging (or CoMA), attains high accuracy on the current task by leveraging plasticity, while not deviating too far from the previous weight configuration, ensuring stability. We also propose an improved variant of CoMA, named Continual Fisher-weighted Model Averaging (or CoFiMA), that selectively weighs each parameter in the weights ensemble by leveraging the Fisher information of the weights of the model. Both variants are conceptually simple, easy to implement, and effective in attaining state-of-the-art performance on several standard CL benchmarks. Code is available at: https://github.com/IemProg/CoFiMA.
Keywords:
Continual Learning Model Averaging.
@@ 1 Introduction
Continually learning from a sequence of tasks with a unified model is a challenging problem due to catastrophic forgetting (CF) [15] – a phenomenon that is marked by deterioration of performance on previously seen data. Continual learning (CL) has emerged as a solution to CF that allows models to assimilate information from new tasks while retaining classification capability for the previously learned classes [36]. Until recently, CL approaches predominantly focused on relatively small networks, often ResNets [18], starting from random initialization [36, 75]. Lately, the prominence of large Pre-Trained Models (PTMs) [11, 27, 51] – Vision Transformer (ViT) [27, 35] pre-trained on large datasets (e.g. ImageNet [55], LAION-400M [58]) – has led to an influx of CL methods that are leveraging the strong representation of the PTMs, causing a paradigm shift in CL [77, 78, 76, 69, 87]. In detail, numerous PTM-based CL methods [81, 10, 40] have empirically validated that a good initial representation, obtained with the help of large and diverse pre-training, facilitates incremental learning since new tasks can be learned with fewer training steps. However, sequential full fine-tuning of the PTM backbone results in deterioration of the original PTM representation, alongside significant forgetting on the previously learned tasks [85, 38, 87, 47]. To counteract overfitting, many methods have been proposed that either heuristically confine the PTM fine-tuning only to the first adaptation session [38, 87, 47] or carefully choose a low learning rate to fine-tune the backbone [85]. Nevertheless, achieving a satisfactory balance between quickly accruing knowledge on new tasks while preserving the generalizability of the PTMs remains an open research question.
Figure 1: Comparison of existing model averaging techniques with our proposed technique for CL. (a) Averaging the weights of the pre-trained model and the one fine-tuned leads to simultaneous improvement in out-of-distribution and target dataset performance. (b) Models soups combines multiple fine-tuned models resulting in a robust unified model. (c) In the proposed CoFiMA the weights of the current and past models are weighted based on their Fisher Information matrices (represented by $\mathcal{F}$), resulting in balanced performance for both the current and old tasks.
In the quest for achieving robust fine-tuning of PTMs, several studies have investigated the application of weight averaging (WA) methods [42, 25, 1, 80, 79, 37]. The essence of these methods is to ensemble several fine-tuned PTMs to obtain a single model that encapsulates the representational capabilities of multiple models. While, to our knowledge, WA has not yet been investigated for CL, two recent studies [80, 79] have sparked our interest in examining the feasibility of WA for CL. Firstly, WiSE-FT [80] improves the robustness of the fine-tuning procedure of zero-shot classifiers like CLIP by averaging the weights of both fine-tuned and pre-trained models (Fig. 1(a)). Here, the resulting ensemble exhibits high accuracy on the target distribution while preserving the out-of-distribution (OOD) performance of the original PTM. It is important to recognize that model robustness and CL intersect: the balance between OOD performance and target performance mirrors the CL stability-plasticity trade-off, where the goal is to accommodate new tasks while preserving efficacy on prior tasks [42, 25, 1]. Secondly, “Model soups” [79] shows that combining multiple fine-tunings of the same PTM through WA enhances performance on both in- and out-of-distribution tasks (Fig. 1(b)). Their experiments underscore the potential of WA with a relatively large pool of models (e.g. 32), mirroring the number of tasks typically encountered in CL. Nevertheless, Model soups do not establish the feasibility of using WA when models are trained on different tasks, a capability required for CL.
Motivated by these observations, we cast PTM-based CL as a robust fine-tuning problem and propose Continual Model Averaging (CoMA) as a response to the PTM fine-tuning conundrum in CL. In CoMA we linearly combine the weights from the previous task ${}{t\!-\!1}$ with the weights of the fully fine-tuned model from the current task $t$ (Fig. 1(c)). Full fine-tuning on the current task ensures plasticity, while ensembling the weights with the previous task checkpoint preserves performance on the previous task. The ensemble at task $t$ is then used as an initialization for the next task ${}{t\!+\!1}$. In essence, our proposed CoMA exploits the model averaging techniques [80, 79] to the CL scenario by extending the model averaging to a sequence of downstream tasks. Intuitively, CoMA is effective because task-specific models optimized sequentially could lie in the same basin of the total loss landscape [41, 43].
@@
The model averaging in CoMA assumes that all the network’s weights have the same importance for a given task. While effective, putting equal importance to all weights could result in a weight-ensembled model that lies in an error basin with high loss [79]. To mitigate this problem, we aim to selectively ensemble the weights based on the importance of each weight parameter for a given task. Inspired by Elastic Weight Consolidation (EWC) [26], and the more recent work of Matena et al. [37], we leverage the Fisher information [9, 62] to weigh the model parameters during model averaging. Fisher information inherently captures the importance of each weight parameter on the dataset (or task) the model has been trained on. We call this variant of CoMA as Continual Fisher-weighted Model Averaging (CoFiMA), and shown in Fig. 1(c). CoFiMA stores only the Fisher values of the previous task, thereby rendering Fisher-weighted averaging compatible with CL constraints. CoFiMA maintains computational efficiency, necessitating only a singular forward and backward pass on the data of the current task to estimate Fisher values [48, 26] and yields state-of-the-art performance surpassing both CoMA and existing PTM-based CL solutions.
Our contributions are summarised as follows:
  • We draw a parallel between robust fine-tuning and continual learning and show that model-averaging is a simple yet effective solution for PTM-based CL problem. We propose CoMA, a weight-ensemble inspired CL approach, that addresses the challenging task of stability-plasticity trade-off.
  • We extend CoMA to CoFiMA by employing Fisher information to adaptively weigh the parameters of the previous and current task model.
  • We run extensive experiments on several standard CL benchmarks and demonstrate that CoFiMA yet being simple, it consistently outperforms PTM-based CL solutions.
@@ 2 Related Work
Continual Learning with PTM. Not a long time ago, the predominant focus in CL has been on the sequential training of deep neural networks from scratch, aiming to proficiently acquire new tasks while mitigating forgetting of preceding tasks. Typical CL strategies encompass regularization-based approaches [26, 2, 84, 33, 8], which maintain the initial model and selectively stabilize parameter or prediction alterations; replay-based approaches [72, 82, 50, 4], that seek to approximate and regenerate previously learned data distributions; and architecture-based approaches [83, 59, 56], which allocate discrete parameter sub-spaces for each task.
@@
Differently, the recent trajectory of CL research has probed into the advantages of PTMs [78, 77, 10]. Representations derived from pre-training have demonstrated the capacity to facilitate not only knowledge transfer but also resilience against catastrophic forgetting during downstream continual learning [52, 40]. Moreover, learning on substantial base classes during the pre-training phase permits CL with minimal adaptations [81]. For example, L2P [78] leveraged techniques inspired by pre-trained knowledge utilization in NLP, employing an additional set of learnable parameters, termed “prompts”, which guide a pre-trained representation layer in learning incremental tasks. DualPrompt [77] elaborated on this concept by attaching supplementary prompts to the pre-trained representation layer to facilitate the learning of both task-invariant and task-specific instructions. Though prompt-based approaches have been documented to significantly outperform conventional CL baselines, they introduce an extra inference cost. Recently, Zhang et al. [85] showed that sequential fine-tuning with a small learning rate using PTMs outperforms traditional CL approaches. Wang et al. [74, 73] propose an architecture that employs multiple narrower sub-networks to manage incremental tasks, effectively reducing generalization errors in CL. However, this approach introduces increased complexity.
@@
Unlike prior approaches in CL that enhance PTMs using prompts [77, 78], ensemble of experts [74, 73, 57] or replay-buffers [72, 82, 50, 4], CoFiMA employs a different strategy. CoFiMA enables plasticity by unlocking all model parameters to be fine-tuned unlike  [38, 78, 77]. Furthermore, it reduces forgetting during training by averaging the weights of parameters from previous models. In contrast to EWC [26], which uses the Fisher Information Matrix (FIM) as a regularization term for L2-transfer between tasks, CoFiMA applies the FIM as a weighting factor to assess the significance of weights for each task, without any regularization constraints (more details in supplementary material).
@@
Output-space/weight-space ensembles. Traditional ensemble methods, or output-space ensembles, combine multiple classifiers’ predictions, often outperforming single models and providing more calibrated uncertainty estimates under distribution shifts [16, 31, 46, 64, 37]. These output-space ensembles, however, demand substantial computational resources at inference. Weight-space ensembles offer a computationally efficient alternative by interpolating between model weights [80, 79, 22, 44, 65]. Wortsman et al. [80] achieved this by interpolating between zero-shot CLIP and fine-tuned model weights, resulting in performance gains on both the fine-tuning task and under-distribution shifts. Matena et al. [37] propose an advanced WA technique that uses Fisher-values for different text classification tasks but they do not investigate the feasibility of this approach for CL. In Federated Learning (FL), model averaging, notably FedAvg, is a fundamental technique for amalgamating insights from decentralized data while upholding privacy [39]. This approach involves training local models on distributed nodes and averaging their parameters to update a global model, thus enhancing learning efficiency and data privacy [39, 32, 24]. Our approach differs from existing WA approaches [80, 79, 37] as it focuses on sequential finetuning with weight-averaging to be accustomed to the CL setting, we iteratively perform our procedure once per task, using the averaged model at each task as the initialization for the next task.
@@
WA is also strongly related to the notion of linear model connectivity, introduced by Frankle et al. [14]. This concept identifies a condition where the accuracy levels are maintained throughout the linear interpolation between the weights of two separate networks. Interpolation of neural network weights has been shown to maintain high accuracy across various scenarios, along a shared optimization trajectory [14, 80, 12, 7, 22, 79]. Analogously, Neyshabur et al. [43] demonstrates that there exists a connection between minima obtained by pre-trained models versus freshly initialized ones. They note that there is no performance barrier between solutions coming from pre-trained models, but there can be a barrier between solutions of different randomly initialized models. Mizradeh et al. [41] investigated linear connectivity in the context of multi-task learning and CL. They show that there exists a linear path solution between two models trained on two tasks “A” and “B”: one excelling in task A, and the other fine-tuned on both tasks A and B. These works offer a solid foundation of both theoretical and empirical evidence supporting the efficacy of linear interpolation in model performance. Building on these findings, our work presents a novel solution tailored to CL, an area yet to be explored.
@@ 3.1 Problem formulation and overview
Continual learning with pre-trained models. We consider a classification model $M_{\varphi}(\cdot)\!=\!h(f(\cdot))$, where $f(\cdot)$ is a feature extractor and $h(\cdot)$ is a classification head, both parameterized by a unified set of parameters $\varphi$. The feature extractor $f_{\boldsymbol{\theta}}$ is initialized with parameters $\boldsymbol{\theta}_{0}$ of a PTM and then trained sequentially on a series of incremental tasks, each represented by the corresponding training set $\mathcal{D}_{t}$, for $t\in\{1,\ldots,T\}$. The primary objective is to achieve robust performance across the test sets associated with these tasks. Specifically, for each task $t$, the dataset $\mathcal{D}_{t}$ is defined as $\mathcal{D}_{t}=\bigcup_{c\in C_{t}}\mathcal{D}_{t}^{c}$ where $\mathcal{D}_{t}^{c}={(\mbox{\bf x}_{n}^{c},\boldsymbol{y}_{n}^{c})}_{n=1}^{N_{c}}$, and $C_{t}$ denotes the set of novel classes introduced in task $t$. Here, $N_{c}$ represents the number of training instances for each class $c$, with $(\mbox{\bf x}_{n}^{c},\boldsymbol{y}_{n}^{c})$ denoting the $n$-th training instance and its corresponding label. In Class-Incremental Learning (CIL), evaluation is conducted across all the observed classes without the need for task-index labels [68].
@@
This problem poses two primary challenges: (i) the necessity to adapt the knowledge acquired from the PTM to new tasks; and (ii) the importance of maintaining the model’s comprehensive learning capabilities to avoid forgetting previously acquired knowledge while assimilating new tasks.
Overview. In this work we propose model averaging as an effective solution for PTM-based CIL. As fine-tuning a PTM on a new task causes the weights to deviate from the original PTM and previous task configuration, model averaging avoids forgetting on a previous task and maintains the generalizability of PTM by averaging the weights of the models of the previous task and current task. Each task concludes with inference on all tasks seen so far using the averaged model and using it as an initialization for the next one. We name this approach Continual Model Averaging (CoMA) and will be discussed in detail in Sec. 3.2.
We extend CoMA by introducing Continual Fisher-weighted Model Averaging (CoFiMA), wherein we average two models at any task $t$ based on an additional weighing coefficient that is determined by Fisher information [62, 9]. Fisher information of a given model parameter dictates the importance of that particular weight towards a task. This procedure is detailed in Sec. 3.3.
Figure 2: Illustration of the parameter trajectory with model averaging in the loss landscape. (a) The trajectory where models are re-initialized at the start of each task leading to disparate solutions. (b) Depicts decreasing $\lambda\!=\!1/t$: Uniform Averaging without re-initialization, showing the convergence of model parameters towards a solution that balances between tasks. (c) Recent tasks are given more weight, resulting in a solution that remains close to the latest task’s model while considering previous tasks.
3.2 Continual Model Averaging (CoMA)